The length of a side of the base of a pyramid is 9 cm, and the slant height is 19 cm. Find the surface area of the pyramid.

SA = ? cm^2

Found 2 solutions by Boreal, solver91311:
Answer by Boreal(15235)

SA+area of base plus 4*area of one of the faces
9*9=81 cm^2 for base
area of the face is a triangle with area=(1/2)bh=(1/2)9*19=171/2 cm^2
There are four of these faces, so their combined area is 342 cm^2
Total area is 423 cm^2
It is also the base area + (1/2) perimeter* slant length=81+36*19/2
=81+ 18*19=81+342=423 cm^2

Answer by solver91311(24713)



Since you only provided the measure of one side of the base, it is safe to presume that the base is equilateral. But you failed to mention the shape of the base polygon. Triangle? Square? Decagon? Triacontakaipentagon?

You also failed to mention whether you want the Total Surface Area or the Lateral Surface Area.

Please provide ALL of the information the next time you ask a question. We can't see your book, worksheet, class notes, whatever and we can't smell what you mean.

John

My calculator said it, I believe it, that settles it